OpenStudy (anonymous):

Find the algebraic expression that is equal to . sec(-theta)sin(-theta)cot(-theta)?how do i find it? Answer 1

OpenStudy (anonymous):

can u please work the problem

OpenStudy (anonymous):

how did u get 1 as an answer?

OpenStudy (mertsj):

\[\sec (-\theta)\sin (-\theta)\cot (-\theta)=\sec \theta(-\sin \theta)(-\cot \theta)=\] \[\frac{1}{\cos \theta}(-\sin \theta)(-\frac{\cos \theta}{\sin \theta})=\]

OpenStudy (mertsj):

\[1\]

OpenStudy (anonymous):

thank u

OpenStudy (mertsj):

yw